# Forward and Backward State Abstractions for Off-policy Evaluation

### v1

@@ 4 Numerical experiments
Method. We investigate the finite sample performance of our proposed methods (details in Appendix B), the forward, backward and two-step procedures.
Comparisons. We compare the proposed abstraction obtained via the two-step procedure (denoted by ‘two-step’), single-iteration forward (‘forward’) and backward (‘backward’) abstractions against Markov state abstraction (Allen et al., 2021) (‘Markov’) and a reconstruction-based abstraction (Lange & Riedmiller, 2010) (‘auto-encoder’). Each abstraction’s performance is tested using FQE (Le et al., 2019) applied to the abstract state space. We also report the performance of a baseline FQE applied to the unabstracted, ground state space (‘FQE’).
Environments. We consider two environments from OpenAI Gym (Brockman et al., 2016), “CartPole-v0” and “LunarLander-v2”, with original state dimensions of 4 and 8, respectively. For each environment, we manually include 296 and 292 irrelevant variables in the state, leading to a challenging 300-dimensional system. Refer to Appendix C for more details about these environments.
Results. We report the MSEs and biases of different post-abstraction-OPE estimators and those of the baseline FQE estimator without abstraction in Figure  4 and Figure C.1 in Appendix C. We summarize our findings as follows. First, the proposed two-step method outperforms other baseline methods, with the smallest MSE and absolute bias in all cases. Since ‘Markov’ and ‘auto-encoder’ are types of model-irrelevant abstractions, these comparisons demonstrate the advantages of the proposed two-step method over single-iteration forward and backward procedures. Second, both figures indicate that the baseline FQE applied to the ground state space performs the worst among all cases. This demonstrates the usefulness of state abstractions for OPE.
Figure 4: MSEs and biases of FQE estimators when applied to ground and abstract state spaces with various abstractions. The behavior policy is $\epsilon$-greedy with respect to the target policy, with $\epsilon=0.1,0.3,0.5,0.7$ from left to right.
@@
This appendix is structured as follows: Section A introduces additional related works on confounder selection in causal inference. The implementation details of the proposed state abstraction are discussed in Section B. Additional information concerning the environments and computing resources utilized is presented in Section C. All technical proofs can be found in Section D.
@@ Appendix A Confounder selection in causal inference
Broadly speaking, confounding refers to the problem that even if two variables are not causes of each other, they may exhibit statistical association due to common causes. Controlling for confounding is a central problem in the design of observational studies, and many criteria for confounder selection have been proposed in the literature. A commonly adopted criterion is the “common cause heuristic”, where the user only controls for covariates that are related to both the treatment and the outcome (Glymour et al., 2008; Austin, 2011; Shortreed & Ertefaie, 2017; Koch et al., 2020). Another widely used criterion is to simply use all covariates that are observed before the treatment in time (Rubin, 2009; Hernán & Robins, 2010, 2016). However, both of these approaches are not guaranteed to find a set of covariates that are sufficient to control for confounding. From a graphical perspective, confounder selection is essentially about finding a set of covariates that block all “back-door” paths (Pearl, 2009), but this requires full structural knowledge about the causal relationship between the variables which is often not possible. This motivated some methods that only require partial structural knowledge (Vander Weele & Shpitser, 2011; VanderWeele, 2019; Guo & Zhao, 2023). All the aforementioned methods need substantive knowledge about the treatment, outcome, and covariates. Other methods use statistical tests (usually of conditional independence) to trim a set of covariates that are assumed to control for confounding (Robins, 1997; Greenland et al., 1999; Hernán & Robins, 2010; De Luna et al., 2011; Belloni et al., 2014; Persson et al., 2017). The reader is referred to Guo et al. (2022) for a recent survey of objectives and approaches for confounder selection.
Confounder selection can be considered as a special example of our problem under certain conditions: (i) The state transition is independent, effectively transforming the MDP into a contextual bandit; (ii) The action space is binary, with the target policy consistently assigning either action 0 or action 1, aimed at assessing the average treatment effect; (iii) State abstractions are confined to variable selections. While our proposed two-step procedure shares similar spirits with the aforementioned algorithms, it addresses a more complex problem involving state transitions. Additionally, our focus is on abstraction that facilitates the engineering of new feature vectors, rather than merely selecting a subset of existing ones.
@@ B.1 Implementation details for forward abstraction
The inclusion of the third loss function, $\mathcal{L}_{Q}$, is motivated by the demonstrated benefits of utilizing model-free objectives to guide the training of state abstractions in policy learning, as evidenced by Gelada et al. (2019); Ha & Schmidhuber (2018); François-Lavet et al. (2019). Given our interest in OPE, we integrate the following FQE loss into the objective function,
@@
Environment Hyper-parameters Values Hyper-parameters Values
CartPole-v0 $\alpha_{1}$ 1 $\alpha_{2}$ 1
$\beta_{1}$ 1 $\beta_{2}$ 1
$\gamma_{1}$ 1 $\gamma_{2}$ 1
$\lambda_{1}$ $\min(1,\frac{20}{m})$ $\lambda_{2}$ $\min(1,\frac{10}{m})$
$C_{0}$ 1 $C_{0}$*∗ ∗
$C_{1}$ 1 $C_{1}$ 1
$d_{0}$ $0.15m$ $d_{0}$ $0.15m$
LunarLander-v2 $\alpha_{1}$ 1 $\alpha_{2}$ 1
@@ B.2 Implementation details for backward abstraction
$\alpha_{2}\mathcal{L}_{\rho}+\beta_{2}\mathcal{L}_{ratio}+\delta_{2}\mathcal{L}_{inv}+\lambda_{2}\mathcal{L}_{s},$
@@
Recall that backward-model-irrelevance requires both $\rho^{\pi}$-irrelevance (Definition 6) and (8). The first loss function $\mathcal{L}_{\rho}$ is designed to enforce $\rho^{\pi}$-irrelevance, specified as
$\mathcal{L}_{\rho}=\frac{1}{|\mathcal{D}|}\sum_{(S,A)\in\mathcal{D}}\big{[}\widehat{\rho}^{\pi}(A,S)-\rho^{\pi}_{\phi_{back}}\big{(}A,\phi_{back}(S)\big{)}\big{]}^{2},$
where $\widehat{\rho}^{\pi}$ denotes some consistent estimator of the IS ratio. Note that in two-step procedure, we should replace $\widehat{\rho}^{\pi}(A,S)$ by:
$\widehat{\rho}^{\pi}_{for}(A,\phi_{for}(S))=\frac{\pi_{\phi_{for}}(A|\phi_{for}(S))}{\widehat{b}(A|\phi_{for}(S))}=\frac{\pi(A|S)}{\widehat{b}(A|\phi_{for}(S))},$
where $\widehat{b}$ is estimated from the abstracted experiences and $\pi(A|S)$ keeps static due to the $\pi$-irrelevance property of forward abstraction.
As commented in Section 3.2, the second condition of (8) holds by satisfying the conditional independence assumption between $(A_{t},\phi(S_{t}))$ and $S_{t+1}$ given $\phi(S_{t+1})$. By Bayesian formula, we can show that it is satisfied by the inverse-model-irrelevance and density-ratio-irrelevance when setting the learning policy $\pi$ to $b$. This motivates us to leverage the two objectives $\mathcal{L}_{inv}$ and $\mathcal{L}_{ratio}$ used by Allen et al. (2021) for training MSA. More details regarding these losses can be found in Section 5 of Allen et al. (2021). Note that to obtain non-sequential states $(s,\tilde{s})$ used in $L_{ratio}$, we flip $s^{\prime}$ in the pairs $(s,s^{\prime})$ in each batch instead of shuffling.
@@ C.2 Experimental settings and additional results
For both environments we use Adam Kingma & Ba (2014) optimizer, with learning rate $0.001$ in Cartpole and $0.003$ in LunarLander. Model architectures and hyper-parameters are outlined in B. When conducting OPE, the FQE network has $3$ hidden layers with $64$ nodes per hidden layer for abstraction methods, and is equipped with $5$ hidden layers with $128$ nodes per hidden layer for non-abstracted observations (shown as ‘FQE’ in the plot).
C.2.1 CartPole-v0
Data generating processes
We manually insert 296 irrelevant features in the state, each following a first order auto-regressive model (AR(1))
$\mathbb{P}(S_{t+1,j}|S_{t},A_{t})=\mathbb{P}(S_{t+1,j}|S_{t,j}),~{}~{}~{}~{}~{}j=5,\dots,300.$
We also define a new state-action-dependent reward as
$\mathcal{R}(s_{t},a_{t})=1-2s_{t,1}^{2}-5s_{t,3}^{2},$
where $s_{t,1}$ and $s_{t,3}$ are the first feature (cart position) and third feature (pole angle) of the state $s_{t}$, to replace the original constant rewards. The number of trajectories $n$ in the offline dataset is chosen from $\{5,8,15,30\}$, where each trajectory contains approximately 40 decision points. The target policy is determined by the pole angle: we push the cart to the left if the angle is negative and to the right if it is positive. Namely,
$\pi(s_{t})=\mathbbm{1}(s_{t,3}>0).$
[…]
Model parameters
For the proposed forward and backward models, we set the abstracted state dimension as $100$. For the two-step method, we apply backward abstraction followed by forward abstraction, reducing the dimension from $300\rightarrow 100\rightarrow 6$ for $\epsilon\in\{0.1,0.3\}$. We change the abstracted dimension to $300\rightarrow 100\rightarrow 2$ for $\epsilon\in\{0.5,0.7\}$.
C.2.2 LunarLander-v2
@@
For forward and backward models, we abstract the original state dimension from $300\rightarrow 100$, and for two-step method we reduce dimensions from $300\rightarrow 50\rightarrow 4$, by first using forward model and then backward model.
@@
Figure C.1: MSEs and biases of FQE estimators when applied to ground and abstract state spaces with various abstractions. The behavior policy is $\epsilon$-greedy with respect to the target policy, with $\epsilon=0.1,0.3,0.5$ from left to right.
@@ C.3 Licences for existing assets
We consider two environments from OpenAI Gym (Brockman et al., 2016), “CartPole-v0” and “LunarLander-v2” with the MIT License and Copyright (c) 2016 OpenAI (https://openai.com).
@@ C.4 Computing resources
C.4.1 CartPole-v0
To build Figure 4, we trained 3 abstraction methods and one non-abstraction method on 4 different sizes of data, each with 30 runs, under 4 $\epsilon$ values. Each run takes approximately 1.5 minutes for four methods on an E2-series CPU with 64GB memory on Google Cloud Platform (GCP). It takes about 12 compute hours to complete all the experiments in the figure.
C.4.2 LunarLander-v2
To build Figure C.1, we trained 3 abstraction methods and one non-abstraction method on 3 different sizes of data, each with 30 runs, under 3 $\epsilon$ values. In average, each run takes approximately 4 minutes for four methods on an E2-series CPU with 64GB memory on GCP. It takes about 18 computation hours to complete all the experiments in the figure.
@@ Appendix D Technical proofs
We provide the detailed proofs of our theorems (Theorems 1, 2, 3, 4) in this section.
@@
D.1 Proof of Theorem 1
We prove Theorem 1 in this subsection. We first prove under $Q^{\pi}$-, $\rho^{\pi}$- or $w^{\pi}$-irrelevance, the corresponding methods remain valid when applied to the abstract state space:
  • $Q^{\pi}$-irrelevance. By definition, $Q^{\pi}$ is the expected return given an initial state $S_{1}$ and $A_{1}$. Under $Q^{\pi}$-irrelevance, the Q-function depends on $S_{1}$ only through $\phi(S_{1})$. It follows that $Q^{\pi}$ equals the expected return given $\phi(S_{1})$ and $A_{1}$, the latter being $Q_{\phi}^{\pi}$ – the $Q$-function when restricted to the abstract state space, i.e., $Q_{\phi}^{\pi}(a,\phi(s))=\sum_{t\geq 1}\gamma^{t-1}\mathbb{E}^{\pi}[R_{t}|A_{1}=a,\phi(S_{1})=\phi(s)]$. It follows that
@@ D.1 Proof of Theorem 1
  • $\rho^{\pi}$-irrelevance. We first establish the equivalence between $\rho^{\pi}$ and $\rho^{\pi}_{\phi}$ – the IS ratio defined on the abstract state space. Under $\rho^{\pi}$-irrelevance, $\rho^{\pi}(a,s)$ becomes a constant function of $x=\phi(s)$. Consequently, for any conditional probability mass function (pmf) $f(s|x)$ such that $\sum_{s\in\phi^{-1}(x)}f(s|x)=1$, we have $\rho^{\pi}(a,s)=\sum_{s\in\phi^{-1}(x)}f(s|x)\rho^{\pi}(a,s)$. By setting $f(s|x)$ to the pmf of $S_{t}=s$ given $A_{t}=a$ and $\phi(S)=x$, it follows that
    $\rho^{\pi}(a,s)=\sum_{s\in\phi^{-1}(x)}\mathbb{P}(S_{t}=s|A_{t}=a,\phi(S_{t})=x)\rho^{\pi}(a,s).$ (D.1)
@@
    $\mathbb{P}(S_{t}=s|A_{t}=a,\phi(S_{t})=x)=\frac{\mathbb{P}(A_{t}=a,S_{t}=s|\phi(S_{t})=x)}{\mathbb{P}(A_{t}=a|\phi(S_{t})=x)}.$
    The denominator equals $b_{\phi,t}(a|x)$, the behavior policy when restricted to the abstract state space at time $t$. Notice that this behavior policy can be non-stationary over time, despite that $b$ being time-invariant. As for the numerator, it is straightforward to show that it equals $b(a|s)\mathbb{P}(S_{t}=s|\phi(S_{t})=x)$. This together with (D.1) yields
    $\rho^{\pi}(a,s)=\sum_{s\in\phi^{-1}(x)}\frac{\pi(a|s)}{b_{\phi,t}(a|x)}\mathbb{P}(S_{t}=s|\phi(S_{t})=x)=\frac{\pi_{\phi,t}(a|x)}{b_{\phi,t}(a|x)},$ (D.2)
    where $\pi_{\phi,t}$ denotes the target policy confined on the abstract state space at time $t$. The last term in (D.2) is given by $\rho_{\phi,t}^{\pi}$. Consequently, the cumulative IS ratio $\rho_{1:t}^{\pi}$ is equal to $\prod_{k=1}^{t}\rho_{\phi,k}^{\pi}(A_{k},\phi(S_{k}))$. This in turn yields $\mathbb{E}[f_{2}(\rho^{\pi})]=\mathbb{E}[f_{2}(\rho^{\pi}_{\phi})]$.
  • $w^{\pi}$-irrelevance. Similar to the proof under $\rho^{\pi}$-irrelevance, the key lies in establishing the equivalence between $w^{\pi}(a,s)$ and $w^{\pi}_{\phi}(a,\phi(s))$, the latter being the MIS ratio defined on the abstract state space. Once this has been proven, it is immediate to see that $\mathbb{E}[f_{3}(w^{\pi})]=\mathbb{E}[f_{3}(w^{\pi}_{\phi})]$, so that MIS remains valid when applied to the abstract state space.
@@
    $w^{\pi}(a,s)$ $=$ $(1-\gamma)\sum_{s\in\phi^{-1}(x)}\frac{\sum_{t\geq 1}\gamma^{t-1}p_{t}^{\pi}(s)\pi(a|s)}{p_{\infty}(s)b_{\phi}(a|x)}\mathbb{P}(S=s|\phi(S)=x)$
    $=$ $\frac{(1-\gamma)\sum_{s\in\phi^{-1}(x)}\sum_{t\geq 1}\gamma^{t-1}p_{t}^{\pi}(s)\pi(a|s)}{p_{\infty}(x)b_{\phi}(a|x)}.$
@@
Finally, we establish the validity of DRL. According to the doubly robustness property, DRL is valid when either $Q^{\pi}$ or $w^{\pi}$ is correctly specified. Under $Q^{\pi}$-irrelevance, we have $Q^{\pi}(a,s)=Q^{\pi}_{\phi}(a,\phi(s))$ and thus DRL remains valid when applied to the abstract state space. Similarly, we have $w^{\pi}(a,s)=w^{\pi}_{\phi}(a,\phi(s))$ under $w^{\pi}$-irrelevance, which in turn implies DRL’s validity. This completes the proof.
D.2 Proof of Theorem 2
We prove Theorem 2 in this subsection.
  • For any $s^{(1)}$ and $s^{(2)}$ satisfies (2), we aim to prove
    $Q^{\pi}(a,s^{(1)})=Q^{\pi}(a,s^{(2)}).$
    Toward that end, we use the induction method. Denote

### v2

@@ 5 Numerical experiments
We investigate the finite sample performance of our proposal in this section and detail its implementation in Appendix B.
Comparisons. We compare the proposed deep state abstraction (denoted by ‘DSA’) against single-iteration forward (‘forward’), backward (‘backward’) abstractions, the Markov state abstraction (Allen et al.,, 2021) (‘MSA’) and a reconstruction-based abstraction (Lange and Riedmiller,, 2010) (‘auto-encoder’). For fairer comparison, each abstraction’s performance is tested by applying a base FQE algorithm (Le et al.,, 2019) applied to the abstract state space. We also report the performance of FQE applied to the unabstracted, ground state space (‘FQE’).
Environment. We consider the “LunarLander-v2” environment in this section, with an original state dimension of 8. We manually include 292 irrelevant variables in the state, leading to a challenging 300-dimensional system. Refer to Appendix C for more details about the environment.
Results. We report MSEs and absolute biases of different post-abstraction-OPE estimators and those of the baseline FQE estimator without abstraction in Figure 5. It can be seen that the proposed DSA method outperforms other baseline methods, with the smallest relative MSE and absolute bias in most cases. To conclude, our analysis answers the following questions:
  1. 1.
    Is state abstraction useful for OPE? Both figures show that the baseline FQE applied to the ground state space performs the worst among all cases. This comparison reveals the usefulness of state abstractions for OPE.
  2. 2.
    Is the deep/iterative procedure more effective compared to single-iteration procedure? Notice that ‘Markov’ and ‘auto-encoder’ are types of model-irrelevant abstractions. The comparisons against these abstractions as well as ‘forward’ and ‘backward’ demonstrate the advantages of DSA over single-iteration procedures.
Figure 5: Relative MSEs and absolute biases of FQE estimators when applied to ground and abstract state spaces with various abstractions. The behavior policy is $\epsilon$-greedy with respect to the target policy, with $\epsilon=0.1,0.3,0.5$ from left to right.
@@
This appendix is structured as follows: Section A introduces additional related works on confounder selection in causal inference. The implementation details of the proposed state abstraction are discussed in Section B. Additional information concerning the environment and computing resources utilized is presented in Section C. All technical proofs can be found in Section D.
@@ Appendix A Confounder selection in causal inference
Broadly speaking, confounding refers to the problem that even if two variables are not causes of each other, they may exhibit statistical association due to common causes. Controlling for confounding is a central problem in the design of observational studies, and many criteria for confounder selection have been proposed in the literature. A commonly adopted criterion is the “common cause heuristic”, where the user only controls for covariates that are related to both the treatment and the outcome (Glymour et al.,, 2008; Austin,, 2011; Shortreed and Ertefaie,, 2017; Koch et al.,, 2020). Another widely used criterion is to simply use all covariates that are observed before the treatment in time (Rubin,, 2009; Hernán and Robins,, 2010, 2016). However, both of these approaches are not guaranteed to find a set of covariates that are sufficient to control for confounding. From a graphical perspective, confounder selection is essentially about finding a set of covariates that block all “back-door” paths (Pearl,, 2009), but this requires full structural knowledge about the causal relationship between the variables which is often not possible. This motivated some methods that only require partial structural knowledge (Vander Weele and Shpitser,, 2011; VanderWeele,, 2019; Guo and Zhao,, 2023). All the aforementioned methods need substantive knowledge about the treatment, outcome, and covariates. Other methods use statistical tests (usually of conditional independence) to trim a set of covariates that are assumed to control for confounding (Robins,, 1997; Greenland et al.,, 1999; Hernán and Robins,, 2010; De Luna et al.,, 2011; Belloni et al.,, 2014; Persson et al.,, 2017). The reader is referred to Guo et al., (2022) for a recent survey of objectives and approaches for confounder selection.
Confounder selection can be considered as a special example of our problem under certain conditions: (i) The state transition is independent, effectively transforming the MDP into a contextual bandit; (ii) The action space is binary, with the target policy consistently assigning either action 0 or action 1, aimed at assessing the average treatment effect; (iii) State abstractions are confined to variable selections. While our proposed iterative procedure shares similar spirits with the aforementioned algorithms, it addresses a more complex problem involving state transitions. Additionally, our focus is on abstraction that facilitates the engineering of new feature vectors, rather than merely selecting a subset of existing ones.
@@ B.1 Implementation details for forward abstraction
The inclusion of the third loss function, $\mathcal{L}_{Q}$, is motivated by the demonstrated benefits of utilizing model-free objectives to guide the training of state abstractions in policy learning François-Lavet et al., (2019).
Given our interest in OPE, we integrate the following FQE loss into the objective function,
@@
Hyper-parameters Values Hyper-parameters Values
$\alpha_{1}$ 1 $\alpha_{2}$ 1
@@ B.2 Implementation details for backward abstraction
$\alpha_{2}\mathcal{L}_{b}+\beta_{2}\mathcal{L}_{ratio}+\delta_{2}\mathcal{L}_{inv}+\lambda_{2}\mathcal{L}_{s},$
@@
Recall that backward-model-irrelevance requires both $\rho^{\pi}$-irrelevance (Definition 6) and (5). To enforce $\rho^{\pi}$-irrelevance, we first introduce behavior-irrelevance, this can be achieved by minimizing the following cross-entropy loss for behavior:
$\mathcal{L}_{b}=-\frac{1}{|\mathcal{D}|}\sum_{(S,A)\in\mathcal{D}}\log b\big{(}A=a|S\big{)}$
and followed by concatenating with $\{\pi(a|\bullet):a\in\mathcal{A}\}$, which ensures $\pi$-irrelevance. Note that in deeply-abstracted procedure, we replace the behavior loss by
$\mathcal{L}_{b}=-\frac{1}{|\mathcal{D}|}\sum_{(S_{t},A_{t})\in\mathcal{D}}\log b\big{(}A_{t}=a_{t}|\phi_{1}(S_{t}),\{A_{t-k},\phi_{1}(S_{t-k})\}_{k=1,2,...}\big{)}$
which incorporates history information as mentioned in 4.3. In practice, we do not use all the history information for behaviour policy, instead we use the history up to past two steps: $b(A_{t}|\phi_{1}(S_{t}),A_{t-1},\phi_{1}(S_{t-1}),A_{t-2},\phi_{1}(S_{t-2}))$.
As commented in Section 4.2, (5) holds by satisfying the conditional independence assumption between $(A_{t},\phi(S_{t}))$ and $S_{t+1}$ given $\phi(S_{t+1})$. By Bayesian formula, we can show that it is satisfied by the inverse-model-irrelevance and density-ratio-irrelevance when setting the learning policy $\pi$ to $b$. This motivates us to leverage the two objectives $\mathcal{L}_{inv}$ and $\mathcal{L}_{ratio}$ used by Allen et al., (2021) for training MSA. More details regarding these losses can be found in Section 5 of Allen et al., (2021). Note that to obtain non-sequential states $(s,\tilde{s})$ used in $L_{ratio}$, we flip $s^{\prime}$ in the pairs $(s,s^{\prime})$ in each batch instead of shuffling.
@@ C.2 Experimental settings and additional results
For the environment LunarLander, we use Adam Kingma and Ba, (2014) optimizer with learning rate $0.003$. Model architectures and hyper-parameters are outlined in B. When conducting OPE, the FQE network has $3$ hidden layers with $64$ nodes per hidden layer for abstraction methods, and is equipped with $5$ hidden layers with $128$ nodes per hidden layer for non-abstracted observations (shown as ‘FQE’ in the plot).
@@
For forward and backward models, we abstract the original state dimension from $300\rightarrow 10$, and for DSA method we reduce dimensions from $300\rightarrow 100\rightarrow 50\rightarrow 20\rightarrow 6$, by [backward, forward, backward, forward] order.
@@ C.3 Licences for existing assets
We consider the environment from OpenAI Gym (Brockman et al.,, 2016) “LunarLander-v2” with the MIT License and Copyright (c) 2016 OpenAI (https://openai.com).
@@ C.4 Computing resources
To build Figure 5, we trained 3 abstraction methods and one non-abstraction method on 4 different sizes of data, each with 20 runs, under 3 $\epsilon$ values. In average, each run takes approximately 8 minutes for four methods on an E2-series CPU with 64GB memory on GCP. It takes about 32 compute hours to complete all the experiments in the figure.
@@ Appendix D Technical proofs
We provide the detailed proofs of our theorems (Lemma 1, Theorems 1 & 2) in this section.
@@
An auxiliary lemma. To begin with, we introduce the following lemma which validates the unbiasedness of various OPE estimators under the model-free irrelevance conditions in Definitions 57, whose proof is given in Section D.1.
Lemma D.1 (Unbiasedness under model-free irrelevance conditions)
Under $Q^{\pi}$-, $\rho^{\pi}$- or $w^{\pi}$-irrelevance, the corresponding methods are unbiased when applied to the abstract state space, assuming the oracle Q-function or (M)IS ratio is identifiable from the data:
  • Under $Q^{\pi}$-irrelevance, Q-function-based method (Method 1) remains unbiased, i.e., the Q-function $Q_{\phi}^{\pi}$ defined on the abstract space satisfies $\mathbb{E}[f_{1}(Q^{\pi})]=\mathbb{E}[f_{1}(Q^{\pi}_{\phi})]$;
  • Under $\rho^{\pi}$-irrelevance, SIS (Method 2) remains unbiased, i.e., the IS ratio $\rho_{\phi}^{\pi}$ defined on the abstract state space satisfies $\mathbb{E}[f_{2}(\rho^{\pi})]=\mathbb{E}[f_{2}(\rho^{\pi}_{\phi})]$;
  • Under $w^{\pi}$-irrelevance, MIS (Method 3) remains unbiased, i.e., the MIS ratio $w_{\phi}^{\pi}$ defined on the abstract state space satisfies $\mathbb{E}[f_{3}(w^{\pi})]=\mathbb{E}[f_{3}(w^{\pi}_{\phi})]$.
Moreover, when $\phi$ satisfies either $Q^{\pi}$-irrelevance or $w^{\pi}$-irrelevance, DRL (Method 4) remains unbiased, i.e., $Q_{\phi}^{\pi}$ and $w_{\phi}^{\pi}$ defined on the abstract state space satisfy $\mathbb{E}[f_{4}(Q^{\pi},w^{\pi})]=\mathbb{E}[f_{4}(Q^{\pi}_{\phi},w^{\pi}_{\phi})]$.
Lemma D.1 proves the unbiasedness of the four OPE methods presented in Section 3.3 when applied to the abstract state space, under the corresponding irrelevance conditions. Notably, DRL requires weaker irrelevance conditions compared to the Q-function-based method and MIS, owing to its inherent double robustness property.
D.1 Proof of Lemma D.1
We prove Lemma D.1 in this subsection. We first prove that under $Q^{\pi}$-, $\rho^{\pi}$- or $w^{\pi}$-irrelevance, the corresponding methods remain unbiased when applied to the abstract state space:
  • Unbiasedness under $Q^{\pi}$-irrelevance. By definition, $Q^{\pi}$ is the expected return given an initial state $S_{1}$ and $A_{1}$. Under $Q^{\pi}$-irrelevance, the Q-function depends on $S_{1}$ only through $\phi(S_{1})$. It follows that $Q^{\pi}$ equals the expected return given $\phi(S_{1})$ and $A_{1}$, the latter being $Q_{\phi}^{\pi}$ – the $Q$-function when restricted to the abstract state space, i.e., $Q_{\phi}^{\pi}(a,\phi(s))=\sum_{t\geq 1}\gamma^{t-1}\mathbb{E}^{\pi}[R_{t}|A_{1}=a,\phi(S_{1})=\phi(s)]$. It follows that
@@
  • Unbiasedness under $\rho^{\pi}$-irrelevance. We first establish the equivalence between $\rho^{\pi}$ and $\rho^{\pi}_{\phi}$ – the IS ratio defined on the abstract state space. Under $\rho^{\pi}$-irrelevance, $\rho^{\pi}(a,s)$ becomes a constant function of $x=\phi(s)$. Consequently, for any conditional probability mass function (pmf) $f(\bullet|x)$ such that $\sum_{s^{\prime}\in\phi^{-1}(x)}f(s^{\prime}|x)=1$, we have $\rho^{\pi}(a,s)=\sum_{s^{\prime}\in\phi^{-1}(x)}f(s^{\prime}|x)\rho^{\pi}(a,s^{\prime})$. By setting $f(\bullet|x)$ to the pmf of $S_{t}$ given $A_{t}=a$ and $\phi(S)=x$, it follows that
    $\rho^{\pi}(a,s)=\sum_{s^{\prime}\in\phi^{-1}(x)}\mathbb{P}(S_{t}=s^{\prime}|A_{t}=a,\phi(S_{t})=x)\rho^{\pi}(a,s^{\prime}).$ (D.1)
@@
    $\mathbb{P}(S_{t}=s^{\prime}|A_{t}=a,\phi(S_{t})=x)=\frac{\mathbb{P}(A_{t}=a,S_{t}=s^{\prime}|\phi(S_{t})=x)}{\mathbb{P}(A_{t}=a|\phi(S_{t})=x)}.$
    The denominator equals $b_{\phi,t}(a|x)$, the behavior policy when restricted to the abstract state space at time $t$. Notice that this behavior policy can be non-stationary over time, despite that $b$ being time-invariant. As for the numerator, it is straightforward to show that it equals $b(a|s^{\prime})\mathbb{P}(S_{t}=s^{\prime}|\phi(S_{t})=x)$. This together with (D.1) yields
    $\rho^{\pi}(a,s)=\sum_{s^{\prime}\in\phi^{-1}(x)}\frac{\pi(a|s^{\prime})}{b_{\phi,t}(a|x)}\mathbb{P}(S_{t}=s^{\prime}|\phi(S_{t})=x)=\frac{\pi_{\phi,t}(a|x)}{b_{\phi,t}(a|x)},$ (D.2)
    where $\pi_{\phi,t}$ denotes the target policy confined on the abstract state space at time $t$, namely, $\pi_{\phi,t}(a|x)=\sum_{s^{\prime}\in\phi^{-1}(x)}\pi(a|s^{\prime})\mathbb{P}(S_{t}=s^{\prime}|\phi(S_{t})=x)$. The last term in (D.2) is given by $\rho_{\phi,t}^{\pi}$. Consequently, the sequential IS ratio $\rho_{1:t}^{\pi}$ is equal to $\prod_{k=1}^{t}\rho_{\phi,k}^{\pi}(A_{k},\phi(S_{k}))$. This in turn yields $\mathbb{E}[f_{2}(\rho^{\pi})]=\mathbb{E}[f_{2}(\rho^{\pi}_{\phi})]$.
  • Unbiasedness under $w^{\pi}$-irrelevance. Similar to the proof under $\rho^{\pi}$-irrelevance, the key lies in establishing the equivalence between $w^{\pi}(a,s)$ and $w^{\pi}_{\phi}(a,\phi(s))$, the latter being the MIS ratio defined on the abstract state space. Once this has been proven, it is immediate to see that $\mathbb{E}[f_{3}(w^{\pi})]=\mathbb{E}[f_{3}(w^{\pi}_{\phi})]$, so that the MIS remains Fisher consistent when applied to the abstract state space.
@@
    $w^{\pi}(a,s)$ $=$ $(1-\gamma)\sum_{s^{\prime}\in\phi^{-1}(x)}\frac{\sum_{t\geq 1}\gamma^{t-1}p_{t}^{\pi}(s^{\prime})\pi(a|s^{\prime})}{p_{\infty}(s^{\prime})b_{\phi}(a|x)}\mathbb{P}(S=s^{\prime}|\phi(S)=x)$
    $=$ $\frac{(1-\gamma)\sum_{s^{\prime}\in\phi^{-1}(x)}\sum_{t\geq 1}\gamma^{t-1}p_{t}^{\pi}(s^{\prime})\pi(a|s^{\prime})}{p_{\infty}(x)b_{\phi}(a|x)}.$
@@
Finally, we establish the unbiasedness of DRL. According to the doubly robustness property, DRL is Fisher consistent when either $Q^{\pi}$ or $w^{\pi}$ is correctly specified. Under $Q^{\pi}$-irrelevance, we have $Q^{\pi}(a,s)=Q^{\pi}_{\phi}(a,\phi(s))$ and thus DRL remains Fisher consistent when applied to the abstract state space. Similarly, we have $w^{\pi}(a,s)=w^{\pi}_{\phi}(a,\phi(s))$ under $w^{\pi}$-irrelevance, which in turn implies DRL’s unbiasedness. This completes the proof.
D.2 Proof of Lemma 1
We prove Lemma 1 in this subsection. We first show $Q^{\pi}$-irrelevance under model-irrelevance &\&& $\pi$-irrelevance, and prove the Fisher consistency of the Q-function-based method. Then, we establish Fisher consistency of MIS. Next, we derive Fisher consistency of SIS. Finally, we prove the Fisher consistency of DRL.
  • Fisher consistency of Q-function-based method. We first use the induction method to prove that
    $Q^{\pi}(a,s^{(1)})=Q^{\pi}(a,s^{(2)}),$ (D.3)
    whenever $s^{(1)}$ and $s^{(2)}$ satisfy $\phi(s^{(1)})=\phi(s^{(2)})$. This demonstrates $Q^{\pi}$-irrelevance, which further implies $\mathbb{E}[f_{1}(Q^{\pi})]=\mathbb{E}[f_{1}(Q^{\pi}_{\phi})]$ according to Lemma D.1. We next establish the identifiability of $Q^{\pi}_{\phi}$.
    Define